Find the pHof a 0.600 M HCN solution(Ka= 4.9× 10–10).

(a)3.56

(b)4.77

(c)5.05

(d)6.25

(e)7.61
Answer #1

for weak acids

pH = 1/2 [pKa - log C]

C = 0.600 M

pKa = - log Ka = - log [4.9 x 10-10] = 9.3

pH = 1/2 [9.3 - log 0.6]

pH = 1/2 [9.31 + 0.222]

pH = 4.77

answer = option b = 4.77